The heart of a resting adult pumps blood at a rate of 5.00 Liters/min. 


a.) Convert this to cm3/ s. 


b.) What is this rate in m3/s? 


1
Expert's answer
2021-12-03T12:45:10-0500

a)

"5~\frac l{min}=83.3~\frac {cm^3}s,"

b)

"5~\frac l{min}=0.083~\frac {m^3}s."